OpenStudy (anonymous):

Factor 25x2 + 10x + 1. a.(25x + 1)(x + 1) b.(5x + 1)(5x - 1) c.(5x + 1)²

OpenStudy (callisto):

Use (a+b)^2 = a^2+2ab+b^2 For your question, a=5x , and b=1 25x^2 + 10x + 1 = (5x+1)^2